OpenStudy (anonymous):

Solve V = π r2h for h. Show all work. Help?

OpenStudy (asnaseer):

ok, so you have:\[V=\pi r^2h\]and you want to find h.

OpenStudy (asnaseer):

divide both sides by \(\pi r^2\) in order to isolate h

OpenStudy (anonymous):

k

OpenStudy (asnaseer):

sometimes it helps to think of these in terms of pure numbers. so lets just make up an equation:\[6=3\times2\]in order to isolate 2, we just divide both sides by 3 to get:\[\frac{6}{3}=\frac{\cancel{3}\times2}{\cancel{3}}\]which leads to:\[2=2\]

OpenStudy (asnaseer):

hope that helps illustrate the /process/

OpenStudy (anonymous):

thank you

OpenStudy (asnaseer):

yw :)